# Data Management For Large Language Models: A Survey

### v2

@@ Abstract
Data plays a fundamental role in the training of Large Language Models (LLMs). Effective data management, particularly in the formulation of a well-suited training dataset, holds significance for enhancing model performance and improving training efficiency during pretraining and supervised fine-tuning phases. Despite the considerable importance of data management, the current research community still falls short in providing a systematic analysis of the rationale behind management strategy selection, its consequential effects, methodologies for evaluating curated datasets, and the ongoing pursuit of improved strategies. Consequently, the exploration of data management has attracted more and more attention among the research community. This survey provides a comprehensive overview of current research in data management within both the pretraining and supervised fine-tuning stages of LLMs, covering various noteworthy aspects of data management strategy design: data quantity, data quality, domain/task composition, etc. Looking toward the future, we extrapolate existing challenges and outline promising directions for development in this field. Therefore, this survey serves as a guiding resource for practitioners aspiring to construct powerful LLMs through effective data management practices. The collection of the latest papers is available at https://github.com/ZigeW/data_management_LLM.
Data Management For Large Language Models: A Survey
[…]

### v3

# Data Management For Training Large Language Models: A Survey

@@ Abstract
Data plays a fundamental role in training Large Language Models (LLMs). Efficient data management, particularly in formulating a well-suited training dataset, is significant for enhancing model performance and improving training efficiency during pretraining and supervised fine-tuning stages. Despite the considerable importance of data management, the underlying mechanism of current prominent practices are still unknown. Consequently, the exploration of data management has attracted more and more attention among the research community. This survey aims to provide a comprehensive overview of current research in data management within both the pretraining and supervised fine-tuning stages of LLMs, covering various aspects of data management strategy design. Looking into the future, we extrapolate existing challenges and outline promising directions for development in this field. Therefore, this survey serves as a guiding resource for practitioners aspiring to construct powerful LLMs through efficient data management practices.
Data Management For Training Large Language Models: A Survey
Zige Wang1,2thanks:  Work done during Zige Wang’s internship at Huawei Noah’s Ark Lab.  Wanjun Zhong2thanks:  Corresponding author (zhongwanjun1@huawei.com)  Yufei Wang2  Qi Zhu2  Fei Mi2  Baojun Wang2 Lifeng Shang2Xin Jiang2Qun Liu2 1School of Computer Science, Peking University 2Huawei Noah’s Ark Lab zigewang@stu.pku.edu.cn {zhongwanjun1, wangyufei44, zhuqi41, mifei2, puking.w}@huawei.com {Shang.Lifeng, Jiang.Xin, qun.liu}@huawei.com
@@ 1 Introduction
To construct suitable training datasets, data management is vitally important and challenging in both the pretraining and SFT stages of LLMs, which we define as following:
Data management: the process of organizing a well-suited training dataset with collected data, including the data selection, combination and utilization strategies, and the evaluation of the chosen strategies.